OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 5 to 8. The perimeter of the larger rectangle is 72 inches. What is the perimeter of the smaller rectangle? @sammixboo @DivineSolar @77777jeannie77777

OpenStudy (camerondoherty):

Lets start off like this: is 72 divisible by 8 or by 5?

DivineSolar (divinesolar):

72 x (5/8) = 45

OpenStudy (anonymous):

a 115.2 inches b 45 inches c 40 inches d 16.4 inches

OpenStudy (anonymous):

by 8

DivineSolar (divinesolar):

Or write it like this -> (5/8) (72) = 5 (9) = 45

DivineSolar (divinesolar):

Either way you get 45.

OpenStudy (anonymous):

http://www.webmath.com/k8ratio.html

OpenStudy (camerondoherty):

Ok, so what times 8 equals 72?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

thank you guys!!!!!!!!!!! by the way for your help

DivineSolar (divinesolar):

Even though i have answered it i suppose my help is no longer needed :)! Good luck with the rest sumgirl! :)

OpenStudy (camerondoherty):

Yes, so to increase the ratio, you multiply both sides by 9 5 times 9 equals 45 So 45 is the correct answer ^.^

OpenStudy (camerondoherty):

The ratio end up looking like this 45:72

OpenStudy (anonymous):

;) Good job :) @sumgirl good luck!!!

OpenStudy (anonymous):

thanks

DivineSolar (divinesolar):

Sure thing.